What will happen to the current if the voltage is decreased by one half while the resistance is held constant? State the relationship of voltage and current based on your answer.
1
Expert's answer
2021-10-26T10:48:09-0400

"V=iR"

"I_1=\frac{V_1}{R}"

"I_2=\frac{V_2}{R}"

"\frac{I_1}{I_2}=\frac{V_1}{V_2}"

"V_2=\frac{V_1}{2}"

"\frac{I_1}{I_2}=\frac{V_1}{\frac{V_1}{2}}"

"\frac{I_1}{I_2}=2"

"I_2=\frac{I_1}{2}"